OpenStudy (anonymous):

Nathan invested $2,500 into an account that pays 6% annually. Choose the equation below that gives the value of the account, Vn, after n years. Then, determine the value of the account after 17 years.

OpenStudy (anonymous):

A) Vn = 2,500·(1.06)n-1 ; V17 = $6,350.88 B) Vn = 2,500·(1.06)n ; V17 = $6,350.88 C) Vn = 2,500·(1.06)n ; V17 = $6,731.93 D) Vn = 2,500·(1.06)n-1 ; V17 = $6,731.93